The normal boiling point of liquid methyl acetate is 331 K. Assuming that its molar heat of vaporization is constant at 30.7 kJ/mol, the boiling point of CH3COOCH3 when the external pressure is 0.794 atm is K.
Answer #1

Hvap = 30.7 kJ / mol

R = 8.314 x 10^-3 kJ / mol K

T2 = 331 K

P2 = 1.0 atm

T1 = ?

P1 = 0.794 atm

ln (P2 / P1) = Hvap / R [1 / T1 - 1 / T2]

ln (1/0.794) = 30.7 / (8.314 x 10^-3) [1 / T1 - 1 / 331 ]

6.246 x 10^-5 = 1/ T1 - 1/ 331

T1 = 324 K

temperature = 324 K